Mary has 12 more apples than peaches. If she divides each fruit in half she will have 32 pieces of fruit. how many peaches does Mary have?
Answer by Earlsdon(6294)

Let P = number of peaches and A = number of apples. From the problem description, you can write:
A = P+12 and, when Mary divides each fruit in half, she will double the number of pieces of fruit, right?:
2A+2P = 32 Simplift this by dividing both sides by 2.
A+P = 16 Now substitute A = P+12 into this equation to find P, the number of peaches.
(P+12) + P = 16 Simplify this.
2P + 12 = 16 Subtract 12 from both sides.
2P = 4 Divide both sides by 2.
P = 2
Mary has 2 peaches.